OpenStudy (anonymous):

explain how to find the area of a segment of a circle and a sector of a circle?

ganeshie8 (ganeshie8):

Perpendicular Bisector of Chord Passes Through Center you familiar with this proof ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

Okay. to find sector area : 1) find the area of circle 2) multiply the area with \(\frac{\theta}{360}\) \(\theta \textbf{ is the angle formed by sector}\)

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

thats how you find the sector area?

ganeshie8 (ganeshie8):

yh A full circle makes 360 degrees : Area = \(\pi r^2\) A sector makes \(\theta\) degrees : Area = \(\frac{\theta}{360} . \pi r^2\)

ganeshie8 (ganeshie8):

if the sector forms 180 degrees (half circle), then \(\theta = 180\) can you tell what area of sector will be

OpenStudy (anonymous):

okay, so you find the area of the circle , and then times it with \[\theta/360\]

ganeshie8 (ganeshie8):

yup and that makes sense right ? can you answer the question i asked as well... im not sure whether u get it or not thats why lol